OpenStudy (anonymous):

Evaluate 3(a 4b) when a = 2 and b = 5 -66 54 39 -54

OpenStudy (compassionate):

You forgot your middle sign, so I'll place it for you. 3(2 + 4(5)) = 3(22) = 66 3(2 - 4(5)) = 3(-18) = -54 3(2 * 4(5)) = 3(40) = 120 3(2 / 4(5)) = 3(10) = 30